If fie=2x+2, evaluate volume integral where R is the region of the cube bounded by the plane x=0 ,x=1, y=0, y=1 ,z=0, z=1


1
Expert's answer
2021-12-22T03:32:43-0500
"\displaystyle\int_{0}^1\displaystyle\int_{0}^1\displaystyle\int_{0}^1(2x+2)dzdxdy"

"=\displaystyle\int_{0}^1\displaystyle\int_{0}^1(2x+2)dxdy"

"=\displaystyle\int_{0}^1[x^2+2x]\begin{matrix}
 1 \\
 0
\end{matrix}dy"

"=\displaystyle\int_{0}^13dy=3"